OpenStudy (anonymous):

Which expression is the simplified form of the square root of m to the eighth power? so i think i got the answer but not sure if it is correct can anyone tell me if my answer is correct or not and what is the correct answer?

OpenStudy (anonymous):

m^4

OpenStudy (anonymous):

how did you get that i got m^2

OpenStudy (anonymous):

2^8=256, the square root of 258 =16 2^4=16

OpenStudy (anonymous):

so a square root is just m^1/2 power so if you have the square root of m^16 you just need to multiply the two exponents together (1/2)*(16)=8

OpenStudy (anonymous):

but 2^ is 256 not 258

OpenStudy (anonymous):

@Moose23 is it ok if you help me in other questions?

OpenStudy (anonymous):

Yea, sorry that 258 should have been a 256

OpenStudy (anonymous):

ok so will you help me with my other questions?

OpenStudy (anonymous):

I can point you in the right direction

OpenStudy (anonymous):

ok thnx